OpenStudy (anonymous):

Let x(t) and y(t) be two orthogonal signals so that integral x(t)y(t) by dt with (-inf

OpenStudy (anonymous):

\[\int\limits\limits_{-\infty}^{\infty}\]

OpenStudy (nikvist):

\[E_x=\int\limits_{-\infty}^{-\infty}x^2(t)dt\quad,\quad E_y=\int\limits_{-\infty}^{-\infty}y^2(t)dt\] \[E_{x\pm y}=\int\limits_{-\infty}^{-\infty}(x(t)\pm y(t))^2dt=\] \[=\int\limits_{-\infty}^{-\infty}x^2(t)dt+\int\limits_{-\infty}^{-\infty}y^2(t)dt\pm 2\underbrace{\int\limits_{-\infty}^{-\infty}x(t)y(t)dt}_{=0}=\] \[=E_x+E_y\]

OpenStudy (anonymous):

Thanks pal